Help me please :)

Consider two parallel plate capacitors.   The plates on Capacitor B have twice the area as the plates on Capacitor A, but the plates in both Capacitor B and Capacitor A are separated by 100~ \mu \text{m}100 μm of air. If Capacitor A has a capacitance of C_A = 30.4 nFC​A​​=30.4nF, what is the capacitance of Capacitor B?

Show work please!
Answer #1

Capacitance is C=e0*A/d.

Here e0 is constant. So, Ca=e0*A/d and Cb= e0*2A/d

So, Cb= (e0*A/d) *2 = 60.8nF

If unclear then ask, upvote to appreciate.